jigglypuff314 (jigglypuff314):

\[\large \int\limits \frac{ e ^{2x}+2e^x+1 }{ e^x }dx\] I think I just need help with figuring out the u for the u sub for this one :3

jigglypuff314 (jigglypuff314):

I did this for my second step\[\large \int\limits \frac{ (e^x+1)^2 }{ e^x }dx\]

jigglypuff314 (jigglypuff314):

or should I try to separate it?

OpenStudy (anonymous):

Hmm, I was thinking \(u=e^x\) would be nice.

OpenStudy (anonymous):

Alternatively, this doesn't really need a \(u\) sub.

jigglypuff314 (jigglypuff314):

or something like this? \[\int\limits \frac{ e^x(e^x+2) }{ e^x }dx + \int\limits \frac{1}{e^x}dx\]

OpenStudy (anonymous):

\[\begin{split} \int\limits \frac{ e ^{2x}+2e^x+1 }{ e^x }dx &= \int\limits \frac{e ^{2x}}{e^x}+\frac{2e^x}{e^x}+\frac{1 }{ e^x }dx \\ &= \int\limits e^x+2+e^{-x}~dx \\ &= \int\limits e^x~dx+\int\limits2~dx+\int\limits e^{-x}~dx \\ \end{split} \]

jigglypuff314 (jigglypuff314):

ohhh I see :) thanks!

jigglypuff314 (jigglypuff314):

so it would be e^x + 2x - e^(-x) + C ?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

As a note \(e^x-e^{-x} = 2\sinh(x)\). Though it probably doesn't matter here.

jigglypuff314 (jigglypuff314):

yeah, it doesn't, but thanks! :D